# Limits to Predicting Online Speech Using Large Language Models

### v2

@@ Abstract
We study the predictability of online speech on social media, and whether predictability improves with information outside a user’s own posts. Recent theoretical results suggest that posts from a user’s social circle are as predictive of the user’s future posts as that of the user’s past posts. Motivated by the success of large language models, we empirically test this hypothesis. We define predictability as a measure of the model’s uncertainty, i.e., its negative log-likelihood on future tokens given context. As the basis of our study, we collect 10M tweets for “tweet-tuning” base models and a further 6.25M posts from more than five thousand X (previously Twitter) users and their peers. Across four large language models ranging in size from 1.5 billion to 70 billion parameters, we find that predicting a user’s posts from their peers’ posts performs poorly. Moreover, the value of the user’s own posts for prediction is consistently higher than that of their peers’. We extend our investigation with a detailed analysis on what’s learned in-context and the robustness of our findings. From context, base models learn to correctly predict @-mentions and hashtags. Moreover, our results replicate if instead of prompting the model with additional context, we finetune on it. Across the board, we find that predicting the posts of individual users remains hard.
@@
(c) Average bits per character (BPC) required to predict user tweets, with 95% confidence intervals ($N=5102$). Lower values mean greater predictability.
Figure 2: Predictability of a user’s tweets using LLMs. Bits per character (BPC) measures, on average, how many bits are required to predict the next character. Predictability improves with additional context to the model: (i) past user tweets (user context, Fig. 2(b)) (ii) past tweets from the user’s peers (peer context) and (iii) past tweets from random users (control). We plot the average BPC over users in Fig. 2(c). Most of the predictive information is found in the user context, followed by peer and random context. Our results are robust across models with different parameter sizes and tokenizers.
@@ 1 Introduction
Prediction is of fundamental importance for social research Salganik (2019); Salganik et al. (2020). The predictability of different social variables can provide a scientific window on a diverse set of topics, such as, emotion contagion Kramer et al. (2014) and social influence Bagrow et al. (2019); Cristali and Veitch (2022); Qiu et al. (2018), privacy concerns Garcia (2017); Garcia et al. (2018); Li et al. (2012); Jurgens et al. (2017), the behavior of individuals and groups Tyshchuk and Wallace (2018); Nwala et al. (2023), the heterogeneity of networks Colleoni et al. (2014); Aiello et al. (2012), information diffusion Chen et al. (2019); Weng et al. (2014); Bourigault et al. (2014); Guille and Hacid (2012) and more. Of particular importance for the study of digital platforms is the case of online speech. Studies about the predictability of text, however, have long been limited by the weaknesses of available language models.
We revisit the problem of predicting online speech in light of dramatic advances in language modeling. We focus on the central question: How predictable is our online speech using large language models? Such predictive capabilities could inform research on substantial risks – such as user profiling, impersonation and exerting influence on real users Carroll et al. (2023); Weidinger et al. (2021). Inspired by work exploring the possibility of user profiling through their peers Bagrow et al. (2019), we ask the following: How predictable is a social media post given posts from the author’s peers? We contrast the answer with how predictable social media posts are from the authors’ own posts. Through experiments spanning millions of tweets and thousands of subjects, our study provides a detailed picture of the current state of predicting online speech.
@@ 1.1 Contributions
We investigate the predictability of online speech on the social media platform X (Twitter) using a corpus of 6.25M posts (tweets) of 5000 subjects and their peers. An additional 10M tweets is reserved for tweet-tuning models. We use four large language models of increasing size, GPT-2-XL-1.5B, Llama-3-8B, Falcon-40B, and Llama-2-70B. We use these models to estimate the predictability of our subjects’ posts under various settings.
Predictability improves significantly with model size, as expected. However, even the best model performs poorly without additional context. For illustration, most users’ posts are less predictable than posts from financial news accounts (Fig. 9). We therefore test how much prediction benefits from three additional sources of context, specifically:
  1. 1.
    posts from randomly selected users (random context),
  2. 2.
    posts from the user’s peers (peer context),
  3. 3.
    posts from the user themselves (user context).
Contrary to what prior work suggested Bagrow et al. (2019), our findings on state-of-the-art language models suggest that a user’s own posts have significantly more predictive information than posts from their close social ties. Prediction benefits most from user context, followed by peer context, in turn followed by random context. We find that all of them improve predictability significantly, with a large effect size. Our prompting experiments in Figure 2 illustrate these findings.
To understand what is being learned in-context, we further investigate the locations of greatest uncertainty within a post. We find that base models learn to correctly predict @-mentions and hashtags. Consequently, removing both generally improves predictability across the board. The relative comparisons, however, stay unchanged.
Large language models can be sensitive to prompting strategies. To test the robustness of our findings, we replace prompting with finetuning. Specifically, instead of prompting a tweet-tuned model with a specific source of context, we finetune the model on the same source of context. We find that the relative comparisons between user, peer, and random context, remain the same.
To summarize, the extent to which we can predict online speech is limited even with state-of-the-art language models. Our observations do not suggest that peers exert an outsize influence on an individual’s online posts. Concerns that large language models have made our individual expression predictable are not supported by our findings.
@@ 2 Related work
Modeling online speech
Understanding and modeling language usage on social media Bashlovkina et al. (2023); Kern et al. (2016); Schwartz et al. (2013) has become of particular interest to the research community, where Twitter is one of the most studied social media platforms Zhang et al. (2023); Qudar and Mago (2020). Feeding user generated content to language models could serve both as a mirror and a tool to understand and (potentially) steer users more effectively. For example, LLMs can be used to infer / extract sensitive author information from what they post online Staab et al. (2023). They can also be leveraged to predict users’ political leaning, morality or toxicity Jiang and Ferrara (2023). Without focusing on any particular downstream task, we are interested in how well these models are able to model individual users’ online speech. In other words, how well models are able to learn the distribution of what a user might post.
@@
Predicting attributes of a node from its neighbors is a well-known paradigm in machine learning. In the context of social networks however, it often has problematic implications on privacy since it limits the user’s ability to control what can be inferred about them Garcia (2017); Garcia et al. (2018); Li et al. (2012). Work on the feasibility of "shadow profiles" (predicting attributes of non-users from platform users) have introduced a notion of privacy that is collective Garcia (2017); Garcia et al. (2018). Prediction of sensitive attributes of the user such as age, gender, religion etc. is possible from their peers Jurgens et al. (2017).
Bagrow et al. (2019) go beyond predicting sensitive attributes and look at the predictability of users’ online speech. They investigate the theoretical possibility of peer-based user profiling on Twitter. They look at the information content of tweets using a non-parametric estimator, and derive an upper bound on predictability which shows that 8-9 peers suffice to match the predictive information contained in the user’s own posts. This upper bound on predictability implies that there could exist some predictor which is capable of peer-based user profiling. Our work aims to contextualize their results by empirically testing this hypothesis on concrete predictors. We use transformer-based LLMs, which are currently considered to be the state-of-the-art method for language modeling.
@@ 3 Experimental setup
We replicate the experimental setup of Bagrow et al. (2019) to measure the predictability of users’ online speech using large language models. We define predictability (or rather unpredictability) as a measure of the model’s uncertainty, i.e., its negative log-likelihood, on a specific user’s tweets. We observe how model uncertainty changes given additional sources of context, specifically:
@@
We start by describing our data collection process (Section 3.1) and what models we used (Section 3.2). Then, we go into detail about the implementation of our prompting and finetuning experiments (in Section 3.3 and 4.3, respectively). We don’t share the tweets we collected because of X’s Developer Policy. Our code can be found here: https://github.com/socialfoundations/twitter-predictability
@@ 3.1.1 Sampling stage
Sampling was done by collecting a pool of tweets whose authors would serve as our base for picking our subjects. We collected them using the Twitter Firehose API; which allowed us to collect a 1% sub-sample of real-time tweet activity. This collection period lasted roughly 30 days and was done in early 2023 (from 20. January to 10. February), during which we collected 10M tweets (with $\sim$5M unique authors). We sampled $N=5102$ subjects from this pool ($0.1\%$ of authors) for our experiments. We filtered out users that scored high (above 0.5 on a scale of 0 to 1) on the Bot-O-Meter bot detection API Sayyadiharikandeh et al. (2020) as well as users that had a high retweet ratio. This ensured the selection of authentic users who had a sufficient amount of self-authored tweets for our dataset.
@@ 3.1.2 Timeline collection stage
Similarly to Bagrow et al. (2019), we created a second control group (temporal control). However, because of the similarity of the results, we only report results on the random control group (social control) in the main text. For results including the temporal control, please see Appendix B.5.
@@ 3.2 Models
We used four different model families for our experiments, namely GPT-2 Radford et al. (2019), Falcon Almazrouei et al. (2023) as well as Llama-2 Touvron et al. (2023) and Llama-3 Dubey et al. (2024). These represent models that share a similar transformer-based architecture that have recently become popular due to their impressive generative capabilities. However, they differ in number of parameters (ranging between 1.5B-70B), training corpus and tokenizers. We further differentiate between base and finetuned versions of these models:
@@
For more details on tweet-tuning, please see Appendix A.2. We considered using GPT-3 and GPT-4, however the OpenAI API unfortunately does not allow access to log probabilities for all tokens (which is necessary for our analysis), only to the top-5. We used Huggingface’s transformers library Wolf et al. (2020) to load the models and run our experiments.
@@ 3.3 Prompting Experiments
We also quantify the average effect size of different contexts on predictability relative to each other. We do this by calculating the standardized mean difference (SMD) of the negative log-likelihoods calculated using different sources of context, i.e. the SMD of $\Delta^{c1}_{c2}(u)=\bar{L}_{u}^{c1}-\bar{L}_{u}^{c2}$. Aggregating over all users, this gives us a unit-free quantity to estimate the difference in predictability one context offers over another. Following established practice Cohen (1988), we characterize effect sizes up to $0.2\sigma$ to be small, up to $0.8\sigma$ to be medium and and anything above that to be a large effect size. Larger models (>1.5B parameters) were loaded using 8-bit precision, which has little to no impact on performance Dettmers et al. (2022). For more details on why we use BPC, sensitivity to prompting strategy, context size, etc. please refer to Appendix A.3.
@@ 3.4 Finetuning Experiments
In this section we present a second method for estimating predictability, which has three advantages over our previous prompting approach. First, we can fit the entire context into the model compared to prompting where the maximum size of the input sequence is a limitation. Second, it bypasses a common ailment of LLMs: their sensitivity to prompting strategy. Finally, it avoids any limitations in-context learning might have.
@@ 4 Results
Figure 4: Average effect size of context $c_{2}$ relative to $c_{1}$ on user predictability. Darker green means greater improvement in prediction. Model: Llama-2-70B.
In our experiments, we quantify how additional context influences model uncertainty, i.e. predictability. We used two methods of feeding additional information to our model, with complementary strengths and weaknesses: one prompting based approach which evaluates in-context learning and one finetuning based approach. We start with evaluating how in-context learning performs on base models (Section 4.1), and what they learn from the additional context (Section 4.2). Next, we evaluate tweet-tuned models, where we replicate our main findings by finetuning on different types of contexts (Section 4.3). Finally, we conclude that it is surprisingly hard to predict user tweets using large language models, even with additional context (Section 4.4).
@@ 4.1 Added Context Reduces Uncertainty in Base Models
We start with the results of the prompting approach on base models. In Figure 2, we plot the average BPC calculated over all of our subjects for three different models with varying parameter sizes. The no context case was consistently the most unpredictable with highest BPC. Depending on what context was included, the amount of improvement varied. We show that most of the predictive information is found in the user context, followed by peer and random context. This trend is consistent across all three of our models. For users that are hard to predict, we observe slightly larger improvements in model uncertainty (App. B.4).
We also quantify the effect size each context had on predictability. We plot this for each context pair in Figure 4 on Llama-2. This effect size matrix illustrates our finding that all types of context have a large effect size ($>2.5\sigma$) over having no context at all (last column). It also allows us to quantify the effect size of our main finding: user context offers $1.8\sigma$ improvement over peer context — an even bigger improvement than what peer context offers over random context ($1.3\sigma$). While these effect sizes are most pronounced on Llama-2, we found that these results also translate to our other models (Figure 22 in Appendix).
Figure 7: Average model uncertainty on base vs. tweet-tuned models. First token was left out of the analysis (see Appendix B.6 for more detail). After tweet-tuning, in-context learning on additional context offers little to no improvement.
@@ 4.2 Base Models Learn Syntax From Added Context
Figure 8: Average loss curves with standard error for finetuning experiments on GPT-2-XL-tt. For each subject ($N=5102$) we finetune on the specified context, and compute the cross-entropy loss on $\mathcal{T}_{u}^{\texttt{eval}}$. The plotted averages are computed over the loss curves of 1000 subjects.

### v3

@@ Abstract
Our paper studies the predictability of online speech – that is, how well language models learn to model the distribution of user generated content on X (previously Twitter). We define predictability as a measure of the model’s uncertainty, i.e. its negative log-likelihood. As the basis of our study, we collect 10M tweets for “tweet-tuning” base models and a further 6.25M posts from more than five thousand X (previously Twitter) users and their peers. In our study involving more than 5000 subjects, we find that predicting posts of individual users remains surprisingly hard. Moreover, it matters greatly what context is used: models using the users’ own history significantly outperform models using posts from their social circle. We validate these results across four large language models ranging in size from 1.5 billion to 70 billion parameters. Moreover, our results replicate if instead of prompting the model with additional context, we finetune on it. We follow up with a detailed investigation on what is learned in-context and a demographic analysis. Up to 20% of what is learned in-context is the use of @-mentions and hashtags. Our main results hold across the demographic groups we studied.
@@
(c) Average bits per character (BPC) required to predict user tweets, with 95% confidence intervals ($N=5102$). Lower values mean more predictable.
Figure 2: Predictability of a user’s tweets using LLMs. Bits per character (BPC) measures, on average, how many bits are required to predict the next character. Predictability improves with additional context to the model: (i) past user tweets (user context, Fig. 2(b)) (ii) past tweets from the user’s peers (peer context) and (iii) past tweets from random users (control). We plot the average BPC over users in Fig. 2(c) and the estimated entropy rate of the English language from takahashi_cross_2018 as comparison. Most of the predictive information is found in the user context, followed by peer and random context. Our results are robust across models with different parameter sizes and tokenizers.
@@ 1 Introduction
Prediction is of fundamental importance for social research salganik2019bit; salganik2020measuring. The predictability of different social variables can provide a scientific window on a diverse set of topics, such as, emotion contagion kramer_experimental_2014 and social influence bagrow_information_2019; cristali2022using; qiu2018deepinf, privacy concerns garcia_leaking_2017; garcia_collective_2018; li_towards_2012; ciampaglia_writer_2017, the behavior of individuals and groups tyshchuk2018modeling; nwala2023language, the heterogeneity of networks colleoni2014echo; aiello2012friendship, information diffusion chen2019information; weng2014predicting; bourigault2014learning; guille2012predictive and more. Of particular importance for the study of digital platforms is the case of online speech. Understanding and modeling language usage on social media kern2016gaining; schwartz2013personality has become of particular interest to the research community, where Twitter is one of the most studied social media platforms zhang_twhin-bert_2023; qudar2020tweetbert.
We revisit the problem of predicting online speech in light of dramatic advances in language modeling. We focus on the central question: How predictable is our online speech using large language models? Such predictive capabilities could inform research on substantial risks – such as user profiling, impersonation and exerting influence on real users carroll2023characterizing; weidinger2022taxonomy. Inspired by work exploring the possibility of user profiling through their peers bagrow_information_2019, we ask the following: How predictable is a social media post given posts from the author’s peers? We contrast the answer with how predictable social media posts are from the authors’ own posts. Through experiments spanning millions of tweets and thousands of subjects, our study provides a detailed picture of the current state of predicting online speech and the potential risks that stem from it.
@@ 1.1 Contributions
We investigate the predictability of online speech on the social media platform X (Twitter) using a corpus of 6.25M posts (tweets) of 5000 subjects and their peers. An additional 10M tweets was reserved for tweet-tuning models. We test four large language models of increasing size: GPT-2-XL-1.5B, Llama-3-8B, Falcon-40B, and Llama-2-70B. We use these models to estimate the predictability of our subjects’ posts under various settings. We vary the type of context we provide to our model and observe the effect on model uncertainty. We study the following settings: no context, random context, peer context and user context. We present a detailed analysis on what’s learned in-context and the robustness of our findings. Furthermore, we explore how our results apply to different demographic groups. Our contributions can be summarized as follows:
  • Online speech is surprisingly hard to predict In Section 4.4 we show that most users’ posts are less predictable than posts from financial news accounts. However, even with additional context prediction performs relatively poorly. Only our largest model with additional user context (see Llama-2-70B with user context in Fig. 2) is able to approximate the estimated entropy rate of the English language (1.12 bits). Subject from Nigeria are the least predictable (Section 4.5). Up to 20% of the effect size can be attributed to the model learning to predict hashtags and @-mentions (i.e. syntax) with in-context learning (Section 4.2).
  • Predictability depends largely on context In Section 4.1 show that a user’s own posts have significantly more predictive information than posts from their close social ties. Prediction benefits most from user context, followed by peer context, in turn followed by random context. Our prompting experiments in Figure 2 illustrate these findings, which are robust to model choice and evaluation strategy and even replicate across different demographic groups (Section 4.5). We find that all types of context improve predictability significantly, with a large effect size (Fig. 4). Our finetuning experiments from Section 4.3 suggest that whatever predictive information is inside the peer context, can also be found in the user context.
Our results on the predictability of online speech may inspire research on a wide range of topics. These may include research on influence and information propagation on social networks, social homophily and potential risks. To summarize, the extent to which we can predict online speech is limited even with state-of- the-art language models. Our observations do not suggest that peers exert an outsize influence on an individual’s online posts. Concerns that large language models have made our individual expression predictable are not supported by our findings.
@@ 2 Related work
Modeling online speech using language models
Understanding and modeling language usage on social media bashlovkina_smile_2023; kern2016gaining; schwartz2013personality has become of particular interest to the research community, where Twitter is one of the most studied social media platforms zhang_twhin-bert_2023; qudar2020tweetbert. Many works focus on LLM’s ability to predict singular, sensitive attributes of users based on what they post. Users’ gender, location and relationship status staab_beyond_2023 - even users’ political leaning jiang2023retweet, morality or toxicity jiang2023social are predictable. Our central question is: How predictable is our online speech using large language models? Such predictive capabilities could inform research on substantial risks – such as user profiling, impersonation and exerting influence on real users carroll2023characterizing; weidinger2022taxonomy. We do this through the lens of model uncertainty, which gives us a natural information-theoretic interpretation of our results. We use the cross-entropy of the English language shannon_prediction_1951; takahashi_cross_2018 as a baseline [^1] Of course while keeping in mind that this is an imperfect comparison. Social media language differs from conventional language bashlovkina_smile_2023..
@@
Predicting attributes of a node from its neighbors is a well-known paradigm in machine learning. In the context of social networks however, it often has problematic implications on privacy since it limits the user’s ability to control what can be inferred about them garcia_leaking_2017; garcia_collective_2018; li_towards_2012. Work on the feasibility of "shadow profiles" (predicting attributes of non-users from platform users) have introduced a notion of privacy that is collective garcia_leaking_2017; garcia_collective_2018. Prediction of sensitive attributes of the user such as age, gender, religion etc. is possible from their peers ciampaglia_writer_2017.
bagrow_information_2019 go beyond predicting sensitive attributes and look at the predictability of users’ online speech. They investigate the theoretical possibility of peer-based user profiling on Twitter. They look at the information content of tweets using a non-parametric estimator, and derive an upper bound on predictability which shows that 8-9 peers suffice to match the predictive information contained in the user’s own posts. This upper bound on predictability implies that there could exist some predictor which is capable of peer-based user profiling. Our work aims to contextualize their results by empirically testing this hypothesis on concrete predictors. We test this with 15 peers per subject and use transformer-based LLMs, which are currently considered to be the state-of-the-art method for language modeling.
@@ 3 Experimental setup
We replicate the experimental setup of bagrow_information_2019 to measure the predictability of users’ online speech using large language models. We define predictability (or rather unpredictability) as a measure of the model’s uncertainty, i.e., its negative log-likelihood, on a specific user’s tweets. We observe how model uncertainty changes given additional sources of context, specifically:
@@
We start by describing our data collection process (Section 3.1) and what models we used (Section 3.2). Then, we go into detail about the implementation of our prompting and finetuning experiments (in Section 3.3 and 4.3, respectively). We don’t share the tweets we collected because of X’s Developer Policy. Our code can be found here: https://github.com/socialfoundations/twitter-predictability/
@@ 3.1.1 Sampling stage
Sampling was done by collecting a pool of tweets whose authors would serve as our base for picking our subjects. We collected them using the Twitter Firehose API; which allowed us to collect a 1% sub-sample of real-time tweet activity. This collection period lasted roughly 30 days and was done in early 2023 (from 20. January to 10. February), during which we collected 10M tweets (with $\sim$5M unique authors). We sampled $N=5102$ subjects from this pool ($0.1\%$ of authors) for our experiments. We filtered out users that scored high (above 0.5 on a scale of 0 to 1) on the Bot-O-Meter bot detection API sayyadiharikandeh_detection_2020 as well as users that had a high retweet ratio. This ensured the selection of authentic users who had a sufficient amount of self-authored tweets for our dataset.
@@ 3.1.2 Timeline collection stage
Similarly to bagrow_information_2019, we created a second control group (temporal control). However, because of the similarity of the results, we only report results on the random control group (social control) in the main text. For results including the temporal control, please see Appendix B.5.
@@ 3.2 Models
We used four different model families for our experiments, namely GPT-2 radford2019language, Falcon almazrouei_falcon_2023 as well as Llama-2 touvron_llama_2023 and Llama-3 dubey2024llama. These represent models that share a similar transformer-based architecture that have recently become popular due to their impressive generative capabilities. However, they differ in number of parameters (ranging between 1.5B-70B), training corpus and tokenizers. We further differentiate between base and finetuned versions of these models:
@@
For more details on tweet-tuning, please see Appendix A.2. We considered using GPT-3 and GPT-4, however the OpenAI API unfortunately does not allow access to log probabilities for all tokens (which is necessary for our analysis), only to the top-5. We used Huggingface’s transformers library wolf_transformers_2020 to load the models and run our experiments.
@@ 3.3 Prompting Experiments
We also quantify the average effect size of different contexts on predictability relative to each other. We do this by calculating the standardized mean difference (SMD) of the negative log-likelihoods calculated using different sources of context, i.e. the SMD of $\Delta^{c1}_{c2}(u)=\bar{L}_{u}^{c1}-\bar{L}_{u}^{c2}$. Aggregating over all users, this gives us a unit-free quantity to estimate the difference in model uncertainty one context offers over another. Following established practice cohen_statistical_1988, we characterize effect sizes up to $0.2\sigma$ to be small, up to $0.8\sigma$ to be medium and and anything above that to be a large effect size. Larger models (>1.5B parameters) were loaded using 8-bit precision, which has little to no impact on performance dettmers_llmint8_2022. For more details on why we use BPC, sensitivity to prompting strategy, context size, etc. please refer to Appendix A.3.
@@ 3.4 Finetuning Experiments
In this section we present a second method for estimating model uncertainty, which has three advantages over our previous prompting approach. First, we can fit the entire context into the model compared to prompting where the maximum size of the input sequence is a limitation. Second, it bypasses a common ailment of LLMs: their sensitivity to prompting strategy. Finally, it avoids any limitations in-context learning might have.
@@ 4 Results
Figure 4: Average effect size of context $c_{2}$ relative to $c_{1}$ on model uncertainty. Darker green means greater improvement in model uncertainty (the model becomes less uncertain). For example, user context significantly improves model uncertainty by $3.1\sigma$ over having no context (top right corner). Model: Llama-2-70B.
In our experiments, we quantify how additional context influences model uncertainty, i.e. predictability. We used two methods of feeding additional information to our model, with complementary strengths and weaknesses: one prompting based approach which evaluates in-context learning and one finetuning based approach. We start with evaluating how in-context learning performs on base models (Section 4.1), and what they learn from the additional context (Section 4.2). Next, we evaluate tweet-tuned models, where we replicate our main findings by finetuning on different types of contexts (Section 4.3). We also show that it is surprisingly hard to predict user tweets using large language models, even with additional context (Section 4.4). Finally, we provide some insight into how our results extend to different subgroups within our population (Section 4.5).
@@ 4.1 Added Context Reduces Uncertainty in Base Models
We start with the results of the prompting approach on base models. In Figure 2, we plot the average BPC calculated over all of our subjects for four different models with varying parameter sizes. First token was left out of the analysis (see Appendix B.6 for more detail). The no context case was consistently the most unpredictable with highest BPC. Depending on what context was included, the amount of improvement varied. We show that most of the predictive information is found in the user context, followed by peer and random context. This trend is consistent across all of our models. For users that are hard to predict, we observe slightly larger improvements in model uncertainty (App. B.4).
We also quantify the effect size each context had on predictability. We plot this for each context pair in Figure 4 on Llama-2. This effect size matrix illustrates our finding that all types of context have a large effect size ($2.5\sigma$) over having no context at all (last column). It also allows us to quantify the effect size of our main finding: user context offers $1.8\sigma$ improvement over peer context — an even bigger improvement than what peer context offers over random context ($1.3\sigma$). While these effect sizes are most pronounced on Llama-2, we found that these results also translate to our other models (Figure 23 in Appendix).
@@ 4.2 Base Models Learn Syntax From Added Context
Figure 7: Average model uncertainty on base vs. tweet-tuned models. After tweet-tuning, in-context learning on additional context offers little to no improvement.